A sample of a gas is in a sealed container. The pressure of the gas is 795 torr , and the temperature is 25 ∘C . If the temperature changes to 73 ∘C with no change in volume or amount of gas, what is the new pressure, P2, of the gas inside the container?
Answer #1

Given:

Pi = 795 torr

Ti = 25.0 oC

= (25.0+273) K

= 298 K

Tf = 73.0 oC

= (73.0+273) K

= 346 K

use:

Pi/Ti = Pf/Tf

795 torr / 298.0 K = Pf / 346.0 K

Pf = 923 torr

Answer: 923 torr